Obtain the equation of the regression line Y



on X from the following data :



X 3 5 6 8 9 11



Y 2 3 4 6 5 8

1
Expert's answer
2022-02-07T11:18:26-0500

Solution:

Sum of X = 42

Sum of Y = 28

Mean X = 7

Mean Y = 4.6667

Sum of squares (SSX) = 42

Sum of products (SP) = 30


Regression Equation = ŷ = bX + a


b = SP/SSX = 30/42 = 0.71429


a = MY - bMX = 4.67 - (0.71"\times" 7) = -0.33333


ŷ = 0.71429X - 0.33333